# Multi-Modal Artificial Intelligence of Embryo Grading and Pregnancy Prediction in Assisted Reproductive Technology: A Review

### v1

@@ IV-A Embryo grading
Depending on the type of input data, the embryo grading tasks can be divided into two main categories: grading based on static embryo images and grading based on temporal embryo images. The former is usually evaluated using a single image of an embryo developing to a specific stage, such as day 5. While the latter analyzes image sequences of the embryo at multiple consecutive time points, such as day 1 to day 5, to capture its dynamic development process. As shown in Fig. 3, in the ”Task” column: Figure (a) corresponds to the embryo grading task based on the static cleavage stage image (day 3). Embryos are divided into four grades: A (high), B, C, and D (low). The higher the grade, the more regular the arrangement of cells in the embryo. Figure (b) corresponds to the embryo grading task based on the static blastocyst stage image (day 5). Embryos are classified into two grades: ”Good” (high) and ”Bad” (low), and the higher the grade, the more complete the development of the blastocyst. In addition, the dataset for embryo grading task based on temporal embryo images is shown in Fig. 2(d). The input of the dataset is a series of images, and the labels are classified according to the embryo development level, just like static embryo images.
@@
It is important to note that embryo grading tasks are often closely linked to other related tasks, such as cell counting and embryo segmentation. Embryo development grade is closely related to the number of cells. Therefore, accurate cell counting is essential for embryo grading. At the same time, the embryo segmentation task aims to divide the embryo into different regions, such as the ICM and TE, whose morphological characteristics are of great reference value for assessing the embryo. By analyzing these derived tasks in depth, a more complete understanding of the developmental state of the embryo can be achieved, thereby improving the accuracy and reliability of embryo grading. As shown in Fig. 3, in the ”Task” column: Figure (c) corresponds to the cell counting task; Figure (d) corresponds to the blastocyst segmentation task.
@@
| Wu[29] | 3599, day3 image (static) | 2022 | CNN+Attention | 4-embryo grading | 82.98% |
@@
| Liu[31] | 276801, day2-day6 image (static) | 2025 | Transformer | 16-embryo grading | 73.10% |
| You[32] | 5736, day3 image (static) | 2025 | UNet+GCN | 4-embryo grading | 86.89% |
@@
| Liu[36] | 4997, day5 image (7 focal) | 2024 | ResNet50 | 10-embryo grading | 86.53% |
| Thirumalaraju[37] | 1188, day5 image (static) | 2020 | Xception | 5-embryo grading | 64.95% |
| Paya[38] | 3844, day4 and day5 image (static) | 2022 | ResNet50 | 2-embryo grading | 82.77% |
@@
| Kragh[41] | 4032, day5 image (temporal) | 2019 | CNN+RNN | 6-embryo grading | 67.40% |
| Vaidya[23] | 803, day 2,3 and 5 image (temporal) | 2021 | VGG16+LSTM | 5-embryo grading | 100% |
@@
IV-A1 Static embryo grading
When grading embryos based on static embryo images, the current processing strategy is generally using CNN to extract image features and then grade embryos. Septiandri et al. [26] used ResNet50 [48] to perform a 3-grading task on the 3rd day of embryo development, and the final classification Accuracy (ACC) could reach 91.79. Similarly, Mohamed et al. [27] used VGG16 [49] to perform a 4-grading task on the 5th day of embryonic development, and the final ACC was 90.00. In order to combine the advantages of different CNNs, Wu et al. [28] collected 3601 images on the 3rd day of embryo development of 1800 couples for 4-grading, and extracted the features of DenseNet169 [50], Inception V3 [51], ResNet50 and VGG19 last layer for ensemble classification, respectively. And the final ACC was 74.14. Attention mechanism, as a module to enhance feature extraction, is often used in combination with CNN to improve the model’s attention to key regions during embryonic development. Wu et al. [29] added channel-level attention and point-level attention to CNN, enabling the model to adaptively focus on regions related to the embryo grade. Eventually, the ACC of embryo 4-grading on the test set was 79.15, and the ablation experiments also demonstrated the effectiveness of the attention mechanism.
@@ IV-A1 Static embryo grading
The two-stage model is a modeling strategy often used to handle complex tasks. Its core idea is to divide the overall task into two interconnected sub-stages to achieve more efficient and accurate prediction. In the first stage, the model usually performs the initial screening task, aiming to extract potential key regions from the input; In the second stage, the model conducts fine classification on the candidate results generated in the first stage to further optimize the accuracy rate. You et al. [32] and Guo et al. [33][34] extracted the blastocyst segmentation map and Grad-CAM [54] attention map of embryos in the first stage, respectively, so that the model in the second stage could focus on regions of interest with embryo grading.
@@
In addition, since the microscope can take multi-angle embryo images with different focal lengths, some studies have begun to explore methods for embryo grading after fusion of these multi-angle images [35][36]. This multi-angle image fusion strategy is expected to further improve the accuracy and reliability of embryo grading.
IV-A2 Temporal embryo grading
@@
In addition to basic microscopy, Time-Lapse Microscopy (TLM) is often used to observe embryos. This microscope is capable of taking TLM images, which are a set of temporally sequenced embryo images as shown in Fig. 2(d). TLM can capture the changing process of cells over time, showing the process of cell division, morphological change, movement, development, etc. Compared with the static image, TLM images have an additional time dimension. In order to solve the problem of processing multiple time-series images, Abbasi et al. [39][40] extracted the features of a single image by CNN, and directly fused the features of multiple images by addition or concatenation operation for the subsequent embryo grading task. Apart from these simple and direct methods, Kragh et al. [41], Vaidya et al. [23] and Kalyani et al. [42] extracted single image features through Inception V3, VGG16 or ResNet50, and then integrated the features of TLM images through RNN for embryo internal quality grading.
For the embryo grading task, whether it develops into a blastocyst is an important indicator of embryo quality. Wang et al. [43] directly converted the embryo grading task into a blastocyst classification task, and used a multi-task learning method to train two tasks at the same time: multi-frame blastocyst classification task were assisted by the single-frame classification task. On the multi-frame task model, ResNet was used to extract single frame information and then R(2+1)D network [55] (a 3DCNN method) was used to integrate multi-frame information, and the ACC reached 77.74. Due to the natural advantages of Transformer model in processing temporal data, Fordham et al. [44] directly converted temporal images into tokens and input them into Transformer for blastocyst classification, and finally ACC is 62.50.
The above methods based on static or temporal images realize the automation of embryo quality grading, significantly improve the efficiency and accuracy of embryo evaluation, and the grading effect has surpassed that of professional embryologists, which provides a standardized auxiliary tool for embryologists. The details of these methods are shown in the TABLE I.
IV-A3 Cell counting
Although the embryo grading task can effectively evaluate the quality of the embryo, it requires professional embryologists to label the images. However, the label acquisition of embryo cell counting task is easier, and the number of embryo cells can indirectly represent the quality of the embryo, so there are also a considerable number of studies on embryo cell counting. Malmsten et al. [45] used the improved Inception V3 model to count 8 embryonic cells, and the ACC was as high as 93.90. For the same cell counting task, Nguyen et al. [46] not only considered the spatiality of a single image, but also considered the temporality of multiple continuous embryo development images. By simultaneously inputting a sequence of embryo development images into Transformer, the time information was introduced to help the model carry out more accurate embryo cell 4-counting. When performing embryo grading, exfoliated cells around the embryo can interfere with the judgment of automated grading. Therefore, Zhou et al. [47] extracted the region of interest to eliminate exfoliated cells near the embryo during preprocessing, and used adaptive genetic algorithm to find the best parameters of the network, which completed the two tasks of cell counting and embryo grading simultaneously.
@@ IV-A3 Cell counting
| Saeedi[56] | 221, day5 image (static) | 2017 | K-means+ | TE | 77.30% |
@@
IV-A4 Embryo segmentation
Image segmentation of embryos at the cleavage stage and blastocyst stage aims to accurately identify and quantify different structures and cell compositions during embryonic development, thereby providing a reliable basis for embryo quality assessment in ART. At the blastocyst stage, segmentation mainly focuses on Inner Cell Mass (ICM), TrophEctoderm (TE), BlastCoele (BC) and Zona Pellucida (ZP) to support more detailed blastocyst scoring and PGT biopsy. Saeedi et al. [56] was the early work that proposed automated embryo segmentation, using K-means and watershed algorithm to segment ICM and TE, and the final Dice scores were 83.10 and 77.30, respectively. In addition, this work published relevant dataset, a total of 211 embryo images with segmentation labels, which promoted the development of this field. Based on the previous work, many studies [57][58][59] began to segment ICM and TE by improving UNet [70]. In addition, Chen et al. [60] proposed a segmentation model based on spatial modeling based on the prior knowledge that the ICM is in the center and the TE is in the periphery. In addition to the segmentation of ICM and TE, there are many studies [61][62][63][64][65][66][67] which have included BC, ZP and background (BG) regions into the segmentation criteria.
Although the segmentation of blastocyst stage embryos is the mainstream, there are also some works segmenting cleavage stage embryos. At the cleavage stage, segmentation helps to extract key morphological features such as cell number, symmetry, and cytoplasmic distribution, which are closely related to the developmental potential of the embryo. Jamal et al. [68] used DenseUNet to segment the cells by collecting and labeling images of 327 embryos on the third day of development, and the final Dice score was 96.62. Zhang et al. [69] introduced a SAM-based [71] two-branch segmentation method, in which the semantic branch was used to segment fragments, and the instance branch was used to detect and segment cleavage stage embryos. The final mAP on the embryo cells was 87.40, and the Dice score on the fragments was 69.50.
The embryonic cell counting task and the embryo segmentation task are two key computer vision tasks in the field of ART. The embryonic cell counting task mainly focuses on the images of embryos at the cleavage stage. The embryo segmentation task focuses on embryo images at the blastocyst stage. Although the two are at different stages of embryonic development and have different analytical goals, they are both committed to achieving the automation and standardization of embryo evaluation. The details of these methods are shown in the TABLE I and TABLE II.
@@ IV-A4 Embryo segmentation
| Sawada[74] | 470, day1-day5 image (static) | 2021 | CNN+Attention | AUC=64.20% |
@@
| Abbasi[76] | 130,day3 and day5 image (temporal) | 2023 | CNN+Transformer | ACC=77.50% |
| Borna[77] | 252, day 1,2,3 image (temporal) | 2024 | UNet+CNN | ACC=75.00% |
@@
| Mehta[81] | 333, 11 tabular data | 2024 | AdaBoost | ACC=97.50% |
@@
| Mohammad[83] | 10036, 46 tabular data | 2025 | RIMARC | AUC=92.12% |
| Sergeev[84] | 8732, 19 tabular data | 2024 | DNN | AUC=79.80% |
@@
| Charnpinyo[90] | 1099, day5 image (static)+4 tabular data | 2023 | CNN+MLP | AUC=72.00% |
| Kim[91] | 2555, day5 image (static)+1 tabular data | 2024 | Image segmentation+CNN+MLP | AUC=74.10% |
| Liu[92] | 17580, day5 image (static)+16 tabular data | 2022 | CNN+MLP | AUC=77.00% |
@@
IV-B Pregnancy prediction

### v2

@@ 4.1 Embryo grading
Depending on the type of input data, the embryo grading tasks can be divided into two main categories: grading based on static embryo images and grading based on temporal embryo images. The former is usually evaluated using a single image of an embryo developing to a specific stage, such as day 5. While the latter analyzes image sequences of the embryo at multiple consecutive time points, such as day 1 to day 5, to capture its dynamic development process. As shown in Fig. 3, in the ”Task” column: Fig. 3(a) corresponds to the embryo grading task based on the static cleavage stage image (day 3). Embryos are divided into four grades: A (high), B, C, and D (low). The higher the grade, the more regular the arrangement of cells in the embryo. Fig. 3(b) corresponds to the embryo grading task based on the static blastocyst stage image (day 5). Embryos are classified into two grades: ”Good” (high) and ”Bad” (low), and the higher the grade, the more complete the development of the blastocyst. In addition, the dataset for embryo grading task based on temporal embryo images is shown in Fig. 6(d). The input of the dataset is a series of images, and the labels are classified according to the embryo development level, just like static embryo images.
@@
It is important to note that embryo grading tasks are often closely linked to other related tasks, such as cell counting and embryo segmentation. Embryo development grade is closely related to the number of cells. Therefore, accurate cell counting is essential for embryo grading. At the same time, the embryo segmentation task aims to divide the embryo into different regions, such as the ICM and TE, whose morphological characteristics are of great reference value for assessing the embryo. By analyzing these derived tasks in depth, a more complete understanding of the developmental state of the embryo can be achieved, thereby improving the accuracy and reliability of embryo grading. As shown in Fig. 3, in the ”Task” column: Fig.3 (c) corresponds to the cell counting task; Figure.3 (d) corresponds to the blastocyst segmentation task.
@@
| Wu[37] | 3599, day3 image (static) | 2022 | CNN+Attention | 4-embryo grading | 71.06% |
@@
| Liu[39] | 273438, day2-day6 image (static) | 2025 | Transformer | 16-embryo grading | 73.10% |
| Wicaksono [40] | 1226, day3 image (static) | 2021 | ResNet50+SSL | 3-embryo grading | 80.30% |
| Paya [41] | 3014, day4, day5 image (static) | 2022 | ResNet50+SSL | 2-embryo grading | 87.17% |
| You[42] | 5736, day3 image (static) | 2025 | UNet+GCN | 4-embryo grading | 86.96% |
@@
| Sariniya[44] | 220, day3 image (static) | 2025 | Morphology+EfficientNet | 4-embryo grading | 94.30% |
@@
| Liu[46] | 4997, day5 image (7 focal) | 2024 | ResNet50+Attention | 10-embryo grading | 86.53% |
| Nguyen[47] | 1135, day3 image (static) | 2021 | VGG16 | 2-embryo grading | 75.00% |
| Wu[48] | 8660, day5 image (static) | 2024 | ResNet50 | 6-embryo grading | 94.00% |
| Vermilyea [49] | 8886, day3 image (static) | 2020 | ResNet152 | 5-embryo grading | 64.30% |
| Alkindy[50] | 152, day3 image (static) | 2023 | ResNet+Xception | 4-embryo grading | 98.00% |
| Misaghi[51] | 273438, day2-day6 image (static) | 2024 | EfficientNet | 17-embryo grading | 87.00% |
| Thirumalaraju[52] | 2440, day5 image (static) | 2020 | Ensemble | 5-embryo grading | 64.96% |
| Iza[53] | 249, day5 image (static) | 2024 | ResNet50+Attention | 9-embryo grading | 69.80% |
| Kim[54] | 14989, day3 image (static) | 2024 | ResNet50+Multiscale | 2-embryo grading | 95.89% |
| Sariniya[55] | 200, day3 image (static) | 2025 | Morphology+VGG16 | 2-embryo grading | 88.10% |
| Wang[56] | 41279, day1-day6 image (static) | 2024 | Transformer+SSL | 2-embryo grading | 82.30% |
@@
| Kragh[59] | 4032, day5 image (temporal) | 2019 | InceptionV3+LSTM | 6-embryo grading | 67.40% |
| Vaidya[60] | 803, day 2,3 and 5 image (temporal) | 2021 | VGG16+LSTM | 5-embryo grading | 100.00% |
@@
| Canat[64] | 1909, day1-day5 image (temporal) | 2024 | Transformer | 11-embryo grading | 66.30% |
| Shi[65] | 23027, day1-day5 image (temporal) | 2025 | Transformer(spatial+Temporal) | 9-embryo grading | 63.50% |
| Sun[66] | 2586, day1-day5 image (temporal) | 2025 | Transformer(spatial+Temporal) | 3-embryo grading | 86.06% |
@@
| Zhou[69] | 1500, day2-day5 image (static) | 2025 | FPN+Attention | 8-cell counting | 86.92%(f1score) |
| Dong[70] | 4805, day3 image (static) | 2024 | YOLOv7 | 4-cell counting | 92.40%(mAP50) |
4.1.1 Static embryo grading
When grading embryos based on static embryo images, the current processing strategy was generally using CNN to extract image features and then grade embryos. Septiandri et al. [34] used ResNet50 [71] to perform a 3-grading task on the 3rd day of embryo development, and the final classification Accuracy (ACC) could reach 91.79. Similarly, Mohamed et al. [35] used VGG16 [72] to perform a 4-grading task on the 5th day of embryonic development, and the final ACC was 90.00. In order to combine the advantages of different CNNs, Wu et al. [36] collected 3601 images on the 3rd day of embryo development of 1800 couples for 4-grading, and extracted the features of DenseNet169 [73], Inception V3 [74], ResNet50 and VGG19 last layer for ensemble classification, respectively. And the final ACC was 74.14. Attention mechanism, as a module to enhance feature extraction, was often used in combination with CNN to improve the model’s attention to key regions during embryonic development. Wu et al. [37] added channel-level attention and point-level attention to CNN, enabling the model to adaptively focus on regions related to the embryo grade. Eventually, the ACC of embryo 4-grading on the test set was 71.06, and the ablation experiments also demonstrated the effectiveness of the attention mechanism.
@@ 4.1.1 Static embryo grading
In addition to architectural improvements, [40][56] explored training strategy innovations by introducing contrastive learning into the embryo quality grading task. Their approach leverages a self-supervised learning (SSL) framework [77][78], where discriminative and generalizable feature representations were learned through instance discrimination by constructing positive and negative sample pairs. By eliminating the reliance on manually annotated labels, this method effectively enhanced the performance of downstream grading tasks, demonstrating the potential of self-supervised representation learning in embryo image analysis.
While single-stage models offered simplicity and efficiency, they often struggled to balance accuracy and computational cost in complex tasks. To address this, two-stage models had been widely adopted as a more refined alternative. The first stage performed coarse screening to generate candidate regions, while the second stage refined these candidates through fine-grained classification. You et al. [42] and Guo et al. [43][29] extracted the cell segmentation map and Grad-CAM [79] attention map of embryos in the first stage, respectively, so that the model in the second stage could focus on regions of interest with embryo grading. Sariniya et al. [44] employed two branches, one obtaining features through CNN and the other through morphological analysis, and ultimately fused them for embryo grading.
@@
In addition, since the microscope could take multi-angle embryo images with different focal lengths, some studies had begun to explore methods for embryo grading after fusion of these multi-angle images [45][46]. This multi-angle image fusion strategy was expected to further improve the accuracy and reliability of embryo grading.
The core characteristics and implementation details of the aforementioned studies are summarized in TABLE 2. To broaden the scope and enhance the comprehensiveness of this review, we further incorporate a number of related works [47][48][49][50][51][52][53][54][55][56] that share similar technical methodologies or target comparable application tasks. This enables a more systematic and comparative analysis, facilitating a holistic understanding of the current research landscape.
4.1.2 Temporal embryo grading
@@
In addition to basic microscopy, Time-Lapse Microscopy (TLM) was often used to observe embryos. This microscope was capable of taking TLM images, which were a set of temporally sequenced embryo images as shown in Fig. 6(d). TLM could capture the changing process of cells over time, showing the process of cell division, morphological change, movement, development, etc. Compared with the static image, TLM images had an additional time dimension. In order to solve the problem of processing multiple time-series images, Abbasi et al. [57][58] extracted the features of a single image by CNN, and directly fused the features of multiple images by addition or concatenation operation for the subsequent embryo grading task. Apart from these simple and direct methods, Kragh et al. [59], Vaidya et al. [60] and Kalyani et al. [61] extracted single image features through InceptionV3, VGG16 or ResNet50, and then integrated the features of TLM images through RNN for embryo internal quality grading.
To effectively process three-dimensional time series data, Wang et al. [30] and Barhoun[62] adopted R(2+1)D network [80] (a 3DCNN method) to extract features from dynamic embryo images and applied it to the 2-embryo grading task. Due to the natural advantages of Transformer model in processing temporal data, Fordham et al. [63] and Canat[64] directly converted temporal images into tokens and input them into Transformer for blastocyst classification. To better exploit the spatial structural information within embryo images and the temporal dynamics across image sequences, Shi et al. [65] and Sun et al. [66] proposed Transformer-based architectures for spatial and temporal feature extraction. By leveraging the self-attention mechanism to model long-range dependencies, these approaches effectively capture the spatiotemporal evolution patterns inherent in embryonic development, enabling a more comprehensive understanding of dynamic morphological changes.
The core characteristics and implementation details of the aforementioned studies are summarized in TABLE 2.
4.1.3 Cell counting
Although embryo grading could directly reflect the embryo development potential, it needed professional embryologists to manually label, and the cost was high. Cell counting had become a research hotspot because of its relatively convenient label acquisition and the number of cells can indirectly characterize the quality of embryos. For example, Malmsten et al. [31] achieved high-precision counting of 8-cell embryos based on the improved InceptionV3 model, with an accuracy of 93.90. Nguyen et al. [67] further introduced the time dimension and fused continuous developmental image sequences through the Transformer model, which significantly improved the counting accuracy. Aiming at the interference of exfoliated cells in embryo grading, Zhou et al. [68] proposed to extract regions of interest in the preprocessing stage to eliminate interference, and combined with adaptive genetic algorithm to optimize network parameters, successfully realized the integrated processing of cell counting and embryo grading.
In addition to image classification-based approaches, object detection techniques were also extensively employed for embryonic cell counting. For example, Zhou et al. [69] integrated an attention mechanism with a Feature Pyramid Network [81] to perform detection and enumeration of embryonic cells. Similarly, Dong et al. [70] utilized the YOLOv7 model [82], achieving a mAP@0.5 score of 92.4, which demonstrates excellent performance in this task. The core characteristics and implementation details of the aforementioned studies are summarized in TABLE 2.
@@ 4.1.3 Cell counting
| Saeedi[83] | 211, day5 image (static) | 2017 | K-means+ | TE | 77.30% |
@@
| Liao[98] | 551, day2, day3 image (static) | 2024 | Faster RCNN+Unsupervised | Cell | 88.00% |
4.1.4 Embryo segmentation
In embryo assessment, the task of embryo segmentation was equally important. Image segmentation of embryos at the cleavage stage and blastocyst stage aimed to accurately identify and quantify different structures and cell compositions during embryonic development, thereby providing a reliable basis for embryo quality assessment in ART. At the blastocyst stage, segmentation mainly focused on Inner Cell Mass (ICM), TrophEctoderm (TE), BlastCoele (BC) and Zona Pellucida (ZP) to support more detailed blastocyst grading and PGT biopsy. Saeedi et al. [83] was the early work that proposed automated embryo segmentation, using K-means and watershed algorithm to segment ICM and TE, and the final Dice scores were 83.10 and 77.30, respectively. In addition, this work published relevant dataset, a total of 211 embryo images with segmentation labels, which promoted the development of this field. Based on the previous work, many studies [84][85][86] began to segment ICM and TE by improving UNet [99]. In addition, Chen et al. [87] proposed a segmentation model based on spatial modeling based on the prior knowledge that the ICM is in the center and the TE is in the periphery. In addition to the segmentation of ICM and TE, there are many studies [89][90][91][92][93][32][94] which had included BC, ZP and background (BG) regions into the segmentation criteria.
Although the segmentation of blastocyst stage embryos was the mainstream, there were also some works segmenting cleavage stage embryos. At the cleavage stage, segmentation helped to extract key morphological features such as cell number, symmetry, and cytoplasmic distribution, which are closely related to the developmental potential of the embryo. Jamal et al. [95] used DenseUNet to segment the cells by collecting and labeling images of 327 embryos on the third day of development, and the final Dice score was 96.62. Zhao et al. [96] segmented the ZP, cytoplasm and pronucleus on the first day of embryonic development, and the final comprehensive dice score was 86.90. Zhang et al. [97] introduced a SAM-based [100] two-branch segmentation method, in which the semantic branch was used to segment fragments, and the instance branch was used to detect and segment cleavage stage embryos. The final mAP on the embryo cells was 87.40, and the Dice score on the fragments was 69.50. Liao et al. [98] proposed a serial processing strategy, which first used the Faster R-CNN model to detect the target of embryonic cells, and then further used the unsupervised segmentation method to achieve accurate cell segmentation based on the positioning information provided by the detection results.
The core characteristics and implementation details of the aforementioned studies are summarized in TABLE 3.
@@ 4.1.4 Embryo segmentation
| Sawada[104] | 470, day5 image (static) | 2021 | ResNet50+Attention | AUC=64.20% |
| Rad[105] | 578, day5 image (static) | 2019 | DeepLab+ResNet50 | AUC=70.90% |
@@
| Abbasi[33] | 130,day3 and day5 image (temporal) | 2023 | ResNet50+Transformer | AUC=77.50% |
| Borna[107] | 252, day 1,2,3 image (temporal) | 2024 | UNet+InceptionV3 | AUC=75.00% |
@@
| Boucret[110] | 1580, day1-5 image (temporal) | 2025 | SSL+LSTM+XGBoost | AUC=64.00% |
| Nagaya[111] | 643, day1-5 image (temporal) | 2022 | ResNet50+Attention | AUC=65.40% |
@@
| Mehta[113] | 333, 11 tabular data | 2024 | AdaBoost | AUC=100.0% |
@@
| Mohammad[115] | 10036, 46 tabular data | 2025 | LR | AUC=92.12% |
| Sergeev[116] | 8732, 19 tabular data | 2024 | MLP | AUC=79.80% |
@@
| Borji[121] | 665244, 94 tabular data | 2024 | Transformer | AUC=99.96% |
| Liu[122] | 48514, 38 tabular data | 2025 | CNN | AUC=88.99% |
| Cordeiro[123] | 183, 180 tabular data | 2025 | RF | AUC=99.65% |
| Gao[124] | 13826, 16 tabular data | 2021 | LR | AUC=74.30% |
| Goyal[125] | 495630, 94 tabular data | 2020 | RF | AUC=86.40% |
| Liu[126] | 1857, 48 tabular data | 2023 | LGBM | AUC=63.40% |
| Wen[127] | 949, 20 tabular data | 2022 | XGBoost | AUC=78.70% |
@@
| Charnpinyo[129] | 1099, day5 image (static)+4 tabular data | 2023 | EfficientNet+MLP | AUC=72.00% |
| Salih[130] | 1394, day5 image (static)+17 tabular data | 2025 | ResNet34+MLP | AUC=91.00% |
| Kim[131] | 2555, day5 image (static)+1 tabular data | 2024 | Image segmentation+ResNet50+MLP | AUC=74.10% |
| Liu[132] | 17580, day5 image (static)+16 tabular data | 2023 | EfficientNet+MLP | AUC=77.00% |
@@
| Duval[134] | 9982, day1-day5 image (temporal)+32 tabular data | 2023 | 3DResNet+XGBoost | AUC=72.70% |
4.2 Pregnancy prediction